# SageAttention2: Efficient Attention with Thorough Outlier Smoothing and Per-thread INT4 Quantization

### v3

@@ Abstract
Although quantization for linear layers has been widely used, its application to accelerate the attention process remains limited. To further enhance the efficiency of attention computation compared to SageAttention while maintaining precision, we propose SageAttention2, which utilizes significantly faster 4-bit matrix multiplication (Matmul) alongside additional precision-enhancing techniques. First, we propose to quantize matrixes $(Q,K)$ to INT4 in a hardware-friendly thread-level granularity and quantize matrixes $(\widetilde{P},V)$ to FP8. Second, we propose a method to smooth $Q$, enhancing the accuracy of INT4 $QK^{\top}$. Third, we propose to use an FP32 Matmul buffer for $PV$ to enhance the accuracy of FP8 $\widetilde{P}V$. The operations per second (OPS) of SageAttention2 surpass FlashAttention2 and xformers by about 3x and 5x on RTX4090, respectively. Comprehensive experiments confirm that our approach incurs negligible end-to-end metrics loss across diverse models—including those for large language processing, image generation, and video generation. The codes are available at https://github.com/thu-ml/SageAttention.
footnotetext: * Equal Contribution.
Figure 1: The upper left figure shows the kernel speedup on RTX4090 GPU. The upper right figure shows the end-to-end inference speedup of generating the first token and performance metrics for the needle-in-a-haystack task (gkamradt, 2023) with a sequence length of 100K on Llama3.1 on L20 GPU. The figure below shows two videos generated by CogvideoX (1.5-5B) using FlashAttention2 and SageAttention2 on RTX4090, showing that SageAttention2 accelerates generation 1.8x with no video quality loss.
@@ 1 Introduction
Due to the quadratic time complexity of attention, efficiently implementing it becomes crucial as sequence lengths increase in real-world applications (Jiang et al., 2024). Several strategies have been developed to mitigate computational demands of attention —such as (1) Linear attention methods (Wang et al., 2020; Choromanski et al., 2020; Yu et al., 2022; Katharopoulos et al., 2020) that reduce complexity to $O(N)$ and (2) Sparse attention methods (Liu et al., 2021; Chu et al., 2021; Li et al., ; Xiao et al., 2023b, 2024; Chen et al., 2023; Jiang et al., 2024; Venkataramanan et al., 2023; Gao et al., 2024; Fu et al., 2024) that selectively process parts of the context — these methods are only suitable for a limited range of models and tasks. Widely used attention methods optimize attention by exploiting hardware capacities to enhance the computation speed, such as FlashAttention (Dao et al., 2022), FlashAttention2 (Dao, 2023), FlashAttention3 (Shah et al., 2024), xformers (Lefaudeux et al., 2022), and SageAttention (Zhang et al., 2024). Those works do not omit the computation of attention over parts of the context and achieve impressive speed and precision performance across various models and tasks.
@@
Motivation. For the two matrix multiplication (Matmul) operations in attention: $QK^{\top}$ and $\widetilde{P}V$, SageAttention accelerates them by quantizing the $QK^{\top}$ to INT8 and uses FP16 Matmul with FP16 accumulators for $\widetilde{P}V$. Moreover, to keep the accuracy of attention, SageAttention proposes smoothing $K$ by eliminating its channel-wise outliers. SageAttention achieves 2 $\times$ and 2.7 $\times$ speedup than FlashAttention2 and xformers and is the first quantized attention that incurs negligible end-to-end metrics loss across language, image, and video generation models. However, SageAttention has two weaknesses. (W1) INT8 Matmul achieves only half the speed of INT4. (W2) FP16 Matmul with FP16 accumulators is only compatible with RTX4090 and RTX3090 GPUs. To leverage the faster INT4 tensor cores for $QK^{\top}$ and using a method that can accelerate $\widetilde{P}V$ on a broader range of GPUs, we propose SageAttention2, which quantizes $Q,K$ to INT4 and $\widetilde{P},V$ to FP8.
@@
Challenges. Quantizing $Q,K$ to INT4 and $\widetilde{P},V$ to FP8 presents significant challenges. For example, when only per-tensor quantizing $Q,K$ to INT4, the text-to-video model CogvideoX will generate a completely blurry video (See Figure 2), and Llama2 only achieves a random-guessing-level accuracy of 25% on MMLU. After investigating deeply, we identified three primary challenges: (C1) The numerical range of INT4 is quite limited compared to FP16 or INT8, leading to significant quantization errors when $Q$ and $K$ have some abnormal values. (C2) We discover that the FP32 accumulator designed for FP8 matrix multiplication in the tensor core (mma.f32.f8.f8.f32) is actually FP22, specifically with 1 sign bit, 8 exponent bits, and 13 mantissa bits. This will lead to an accuracy loss of $\widetilde{P}V$.
Figure 2: An example of quantizing Q, K to INT4 from CogvideoX.
@@
Our approach. To address (C1), we first discovered that the per-block quantization of $Q,K$ in SageAttention does not achieve sufficient accuracy for INT4 quantization. Simultaneously, to avoid the extra latency caused by per-token dequantization, where each GPU thread corresponds to multiple quantization scales, we propose an adequately precise quantization method that incurs no additional dequantization overhead. Specifically, we introduce a per-thread quantization approach based on the mapping between the GPU thread and memory layout of matrices as dictated by the PTX mma instructions. This method groups tokens corresponding to the same thread for quantization and dequantization, ensuring that each thread is associated with a single quantization scale. This approach achieves much better accuracy performance than per-block quantization with no additional dequantization overhead. Second, for the significant channel-wise outliers in matrices $Q$ and $K$, we adopt smoothing $K$ in SageAttention and propose an effective method to remove these outliers in $Q$. Specifically, we propose subtracting the average value of the channel dimension of $Q$, referred to as $\overrightarrow{Q}_{m}$. Then, we add $\overrightarrow{Q}_{m}K$ after the $QK^{\top}$ Matmul to ensure the correctness of the attention computation. To address (C2)—the accuracy loss from using a 22-bit accumulator for FP8 Matmul of $\widetilde{P}V$—, we propose to use an FP32 buffer to accumulate the values from the 22-bit accumulator after each block Matmul of $\widetilde{P}V$. This method confines the errors to the block range. Additionally, we design an optional technique to enhance the accuracy of the 22-bit accumulator. Specifically, we could smooth $V$ by subtracting the average value of its channel dimension and adding the subtracted item to the attention output to maintain the correctness.
@@
Performance. Importantly, we offer a high-performance implementation of SageAttention2 on RTX4090 and L20 GPUs. This implementation achieves a peak performance of 481 TOPS on the RTX4090, outperforming FlashAttention2 and xformers by approximately 3x and 5x, respectively. Note that FlashAttention3 is tailored to and can only be used with the Nvidia Hopper architecture, i.e., H100 and H800. We extensively evaluate the end-to-end metrics of state-of-the-art text, image, and video generation models using SageAttention2. Across all models and tasks, SageAttention2 can be directly adopted in a plug-and-play manner with negligible loss in model performance.
@@ 2.1 FlashAttention
The computation of self-attention can be formulated as follows: $S=QK^{\top}/\sqrt{d},~{}P=\sigma(S),~{}O=PV$, where $\sigma(S)_{ij}=\exp(S_{ij})/\sum_{k}\exp(S_{ik})$ is the softmax operation. The matrices $Q$, $K$, and $V$ each have dimensions $N\times d$, while the matrix $S$, $P$ are $N\times N$. While $d$ is typically small, e.g., 64 or 128, $N$ can be millions. Therefore, the $N\times N$ matrices $(S,P)$ are much larger than $(Q,K,V)$, and a naive implementation suffers from the huge amount of global memory IO for $(S,P)$ reads/writes. FlashAttention (Dao, 2023) proposes to tile $Q$, $K$, and $V$ from the token dimension into blocks $\{Q_{i}\},\{K_{i}\},\{V_{i}\}$ with block size of $b_{q}$, $b_{k}$, $b_{v}$, respectively, where $b_{k}$ = $b_{v}$. Then, to avoid the memory IO for $(S,P)$, it uses online softmax (Milakov & Gimelshein, 2018) to progressively compute each block of $O$, i.e., $O_{i}$:
First, for each block of $\{K_{i}\},\{V_{i}\}$, it computes the following equations iteratively:
$S^{j}_{i}=Q_{i}K_{j}^{\top}/\sqrt{d},~{}~{}(m^{j}_{i},\widetilde{P}^{j}_{i})=\tilde{\sigma}(m^{j-1}_{i},S^{j}_{i}),~{}~{}l_{i}^{j}=$ $\exp(m^{j-1}_{i}-m^{j}_{i})l_{i}^{j-1}+\mathrm{rowsum}(\widetilde{P}^{j}_{i}),$
$O^{j}_{i}=\mathrm{diag}\left(\exp(m^{j-1}_{i}-m^{j}_{i})\right)O^{j-1}_{i}$ $+\widetilde{P}^{j}_{i}V_{j}$
Where $m_{i}^{j}$ and $l_{i}^{j}$ are $b_{q}\times 1$ vectors, which are initialized to $\infty$- ∞ and 00 respectively. The $\tilde{\sigma}()$ is an online softmax operator: $m^{j}_{i}=\max\{m^{j-1}_{i},\mathrm{rowmax}(S^{j}_{i})\},~{}\widetilde{P}_{j}^{i}=\exp(S^{j}_{i}-m^{j}_{i})$.
Finally, the output $O_{i}$ can be computed by $O_{i}=\mathrm{diag}(l_{i}^{j})^{-1}O_{i}^{j}$.
@@ 2.2 Quantization
$(\delta_{A},\hat{A})=\psi(A),~{}~{}(\delta_{B},\hat{B})=\psi(B),~{}~{}\hat{C}=\hat{A}\hat{B},$ $~{}~{}C=\psi^{-1}_{\delta_{A}\delta_{B}}(\hat{C})$ (1)
$\psi$ is a quantizer which converts a high-precision (e.g., FP32 or FP16) matrix $A$ to a low-precision format $\hat{A}$ (e.g., INT4 or FP8) with a scale $\delta_{A}$, and $\psi^{-1}$ is a dequantizer to convert back to high-precision. We should have $\psi^{-1}_{\delta_{A}}(\hat{A})\approx A$. The actual matrix multiplication $\hat{A}\hat{B}$ is carried in low precision. In modern GPUs, low-precision matrix multiplication is usually multiple times faster than higher-precision ones. Many quantizers depend on the numerical format and granularity, e.g., how many elements share a common scale factor. For example, an INT4 per-tensor quantizer first computes the scale as the maximum absolute value of the entire tensor, scales the elements to the maximum representable range of INT4 [-7, +7], and then casts to INT4 with rounding: $\hat{A}=\lceil A/\delta_{A}\rfloor,\delta_{A}=\max(|{A}|)/7$. Likewise, per-token quantizer assigns a scale factor for each token of a tensor: $\hat{A}[i,:]=\lceil A[i,:]/\delta_{A}[i]\rfloor,\delta_{A}[i]=\max(|{A[i,:]}|)/7$. Also, per-channel quantizer assigns a scale factor for each channel of the tensor, i.e., along the channel dimension: $A[:,i]=\lceil A[:,i]/\delta_{A}[i]\rfloor,\delta_{A}[i]=\max(|A[:,i]|)/{7}$. The dequantization process is an element-wise scaling. For example, for per-tensor dequantization, $\psi_{\delta_{A}\delta_{B}}^{-1}(\hat{A}\hat{B})=\hat{A}\hat{B}*\delta_{A}*\delta_{B}$.
@@ 2.3 SageAttention
Based on the block tiling in FlashAttention2 (Dao, 2023), SageAttention (Zhang et al., 2024) quantize $Q,K$ to INT8 in a per-block granularity. Moreover, to keep the quantization accuracy, SageAttention proposes to smooth $K$ first:
$K=K$ $-\text{mean}(K)$
$\hat{Q}_{i}=\lceil Q_{i}/\delta_{Q}\rfloor,~{}~{}\delta_{Q}=\max(|{Q_{i}|})/127,$ $~{}~{}\hat{K}_{j}=\lceil K_{j}/\delta_{K}\rfloor,~{}~{}\delta_{K}=\max(|{K_{j}|})/127$
$S_{ij}=$ $Q_{i}K_{j}^{\top}*\delta_{Q}*\delta_{K}$
Where $Q_{i},K_{j}$ are the tilled block in FlashAttention, and $\text{mean}(K)$ is the average value of the channel dimension of $K$. Additionally, SageAttention keeps $\widetilde{P}V$ as FP16 and uses FP16 Matmul with an FP16 accumulator for $\widetilde{P}V$. However, FP16 Matmul with an FP16 accumulator only has a speedup effect on RTX4090 and RTX3090 GPUs.
3 SageAttention-2

### v4

@@ Abstract
Although quantization for linear layers has been widely used, its application to accelerate the attention process remains limited. To further enhance the efficiency of attention computation compared to SageAttention while maintaining precision, we propose SageAttention2, which utilizes significantly faster 4-bit matrix multiplication (Matmul) alongside additional precision-enhancing techniques. First, we propose to quantize matrices $(Q,K)$ to INT4 in a hardware-friendly thread-level granularity and quantize matrices $(\widetilde{P},V)$ to FP8. Second, we propose a method to smooth $Q$, enhancing the accuracy of INT4 $QK^{\top}$. Third, we propose a two-level accumulation strategy for $\widetilde{P}V$ to enhance the accuracy of FP8 $\widetilde{P}V$. The operations per second (OPS) of SageAttention2 surpass FlashAttention2 and xformers by about 3x and 4.5x on RTX4090, respectively. Moreover, SageAttention2 matches the speed of FlashAttention3(fp8) on the Hopper GPUs, while delivering much higher accuracy. Comprehensive experiments confirm that our approach incurs negligible end-to-end metrics loss across diverse models—including those for language, image, and video generation. The code is available at https://github.com/thu-ml/SageAttention
Machine Learning, ICML
Figure 1: The upper left figure shows the kernel speedup on RTX4090 GPU. The upper right figure shows the end-to-end inference speedup of generating the first token and performance metrics for the Needle-in-a-Haystack task (Kamradt, 2023) with a sequence length of 100K on Llama3.1 on L20 GPU. The figure below shows two videos generated by CogvideoX (1.5-5B) using FlashAttention2 and SageAttention2 on RTX4090, showing that SageAttention2 accelerates generation by 1.8x with no video quality loss.
@@ 1 Introduction
Due to the quadratic time complexity of attention, its efficient implementation becomes crucial as sequence lengths increase in real-world applications (Jiang et al., 2024). Several strategies have been developed to mitigate computational demands of attention —such as (1) Linear attention methods (Wang et al., 2020; Choromanski et al., 2021; Yu et al., 2022; Katharopoulos et al., 2020) that reduce complexity to $O(N)$ and (2) Sparse attention methods (Liu et al., 2021; Chu et al., 2021; Li et al., 2022; Xiao et al., 2024b, a; Chen et al., 2024; Jiang et al., 2024; Venkataramanan et al., 2024; Gao et al., 2024; Fu et al., 2024) that selectively process parts of the context — these methods are only suitable for a limited range of models and tasks. The widely used attention methods optimize attention by exploiting hardware capacities, such as FlashAttention (Dao et al., 2022), FlashAttention2 (Dao, 2024), FlashAttention3 (Shah et al., 2024), xformers (Lefaudeux et al., 2022), and SageAttention (Zhang et al., 2025). These works do not omit computations across the entire sequence and achieve impressive speed and accuracy performance across various tasks.
@@
Motivation. For the two matrix multiplication (Matmul) operations in attention: $QK^{\top}$ and $\widetilde{P}V$, SageAttention accelerates them by quantizing the $QK^{\top}$ to INT8 and uses FP16 Matmul with FP16 accumulators for $\widetilde{P}V$. Moreover, to keep the accuracy of attention, SageAttention proposes smoothing $K$ by eliminating its channel-wise outliers. SageAttention achieves 2 $\times$ speedup than FlashAttention2 and is the first quantized attention that incurs negligible end-to-end metrics loss across language, image, and video generation models. However, SageAttention has two weaknesses. (W1) INT8 Matmul achieves only half the speed of INT4. (W2) FP16 Matmul with FP16 accumulators provides a speedup only on RTX 4090 and RTX 3090 GPUs. To leverage the faster INT4 tensor cores for $QK^{\top}$ and using a method that can accelerate $\widetilde{P}V$ on a broader range of GPUs, we propose to quantize $Q,K$ to INT4 and $\widetilde{P},V$ to FP8.
@@
Challenges. Quantizing $Q,K$ to INT4 and $\widetilde{P},V$ to FP8 presents significant challenges. For example, when only per-tensor quantizing $Q,K$ to INT4, the text-to-video model CogvideoX will generate a completely blurry video, and Llama3 only achieves a random-guessing-level accuracy of 25% on MMLU. After investigating deeply, we identified three primary challenges: (C1) The numerical range of INT4 is quite limited compared to FP16 or INT8, leading to significant quantization errors when $Q$ and $K$ have some abnormal values. (C2) We discover that the FP32 accumulator designed for FP8 matrix multiplication in the tensor core (mma.f32.f8.f8.f32) is actually FP22, specifically with 1 sign bit, 8 exponent bits, and 13 mantissa bits. This will lead to an accuracy loss of $\widetilde{P}V$.
@@
Our approach. To address (C1), we first discover that the per-block quantization of $Q,K$ in SageAttention does not achieve sufficient accuracy for INT4 quantization. Simultaneously, to avoid the extra latency caused by per-token dequantization, where each GPU thread corresponds to multiple quantization scales, we propose an adequately precise quantization method that incurs no additional dequantization overhead. Specifically, we introduce a per-thread quantization approach based on the mapping between the GPU threads and memory layout of matrices as dictated by the PTX mma instructions. This method groups tokens corresponding to the same thread for quantization and dequantization, ensuring that each thread is associated with a single quantization scale. This approach achieves much better accuracy performance than per-block quantization with no additional dequantization overhead. Second, for the significant channel-wise outliers in matrices $Q$ and $K$, we adopt smoothing $K$ in SageAttention and further propose an effective method to remove these outliers in $Q$. Specifically, we propose subtracting the average value of the channel dimension of $Q$, referred to as $\overrightarrow{Q}_{m}$. Then, we add $\overrightarrow{Q}_{m}K$ after the $QK^{\top}$ Matmul to ensure the correctness of the attention computation. To address (C2), the accuracy loss from the 22-bit accumulator for FP8 Matmul of $\widetilde{P}V$, we propose a two-level accumulation strategy that uses an FP32 buffer to accumulate the values from the 22-bit accumulator after each block Matmul of $\widetilde{P}V$. This method confines the errors to the block range. Additionally, we design an optional technique to enhance the accuracy of the 22-bit accumulator. Specifically, we could smooth $V$ by subtracting the average value of its channel dimension and adding the subtracted item to the attention output to maintain the correctness.
@@
Performance. Importantly, we offer a high-performance implementation of SageAttention2 on RTX4090 and L20 GPUs. This implementation achieves a peak performance of 481 TOPS on the RTX4090, outperforming FlashAttention2 and xformers by approximately 3x and 4.5x, respectively. Note that FlashAttention3 is tailored to and can only be used with the Nvidia Hopper architecture. Moreover, SageAttention2 matches the speed of FlashAttention3(fp8) on the Hopper GPUs, while delivering much better accuracy. For example, on CogVideoX-1.5 (Yang et al., 2025), Mochi (Team, 2024) and HunyuanVideo (Kong et al., 2024), our method does not compromise end-to-end accuracy, whereas videos generated using FlashAttention3(fp8) often suffer noticable degradation, as visualized in Fig. 8. We extensively evaluate the end-to-end metrics of state-of-the-art text, image, and video generation models using SageAttention2. SageAttention2 can accelerate models in a plug-and-play way with negligible loss in end-to-end metrics.
@@ 2.1 FlashAttention
The attention computation can be formulated as: $S=QK^{\top}/\sqrt{d},~{}P=\sigma(S),~{}O=PV$, where $\sigma(S)_{ij}=\exp(S_{ij})/\sum_{k}\exp(S_{ik})$. The matrices $Q$, $K$, and $V$ each has dimensionality $N\times d$, and $S$, $P$ are $N\times N$. $d$ is typically small, e.g., 64 or 128, and $N$ can be thousands or millions. The time complexity of attention is $O(N^{2})$, primarily due to two matrix multiplications ($QK^{\top}$ and $PV$), both with complexities of $O(N^{2}d)$. FlashAttention (Dao, 2024) is a GPU-friendly attention implementation, which tiles $Q$, $K$, and $V$ from the token dimension into blocks $\{Q_{i}\}_{i=1}^{n_{q}},\{K_{i}\}_{i=q}^{n_{k}},\{V_{i}\}_{i=1}^{n_{v}}$ with block sizes of $b_{q}$, $b_{k}$, $b_{v}$ tokens, respectively, where $n_{q},n_{k},n_{v}$ are the number of tiles, and $b_{k}$ = $b_{v}$. FlashAttention computes the output matrix $O$ in parallel in tiles. Each streaming multiprocessor (SM) computes a block $O_{i}$ (corresponds to a $Q_{i}$) by iteratively loads $K_{j},V_{j}$ for each $j$, and update the output with online softmax (Milakov & Gimelshein, 2018):
$S_{ij}=$ $Q_{i}K_{j}^{\top}/\sqrt{d},~{}~{}(m_{ij},\widetilde{P}_{ij})=\tilde{\sigma}(m_{i,j-1},S_{ij}),$ (1)
$l_{ij}=$ $\exp(m_{i,j-1}-m_{ij})l_{i,j-1}+\mathrm{rowsum}(\widetilde{P}_{ij}),$
$O_{ij}=$ $\mathrm{diag}\left(\exp(m_{i,j-1}-m_{ij})\right)O_{i,j-1}+\widetilde{P}_{ij}V_{j},$
where $m_{ij}$ and $l_{ij}$ are $b_{q}$-dimenalional vectors, initialized with $\infty$- ∞ and 00 respectively. $\tilde{\sigma}(\cdot)$ is an online softmax operator: $m_{ij}=\max\{m_{i,j-1},\mathrm{rowmax}(S_{ij})\},~{}\widetilde{P}_{ij}=\exp(S_{ij}-m_{ij})$. Finally, the output is computed as $O_{i}=\mathrm{diag}(l_{i,n_{q}})^{-1}O_{i,n_{q}}$.
@@ 2.2 Quantization
$(\delta_{A},\hat{A})=\psi(A),~{}~{}(\delta_{B},\hat{B})=\psi(B),~{}~{}C=\psi^{-1}_{\delta_{A}\delta_{B}}(\hat{A}\hat{B})$
$\psi$ is a quantizer which converts a high-precision (e.g., FP32 or FP16) matrix $A$ to a low-precision format $\hat{A}$ (e.g., INT4 or FP8) with a scale $\delta_{A}$, and $\psi^{-1}$ is a dequantizer to convert back to high-precision. We should have $\psi^{-1}_{\delta_{A}}(\hat{A})\approx A$. The actual matrix multiplication $\hat{A}\hat{B}$ is carried out with low precision. In modern GPUs, low-precision matrix multiplication is usually multiple times faster than higher-precision ones. Quantizers differ in numerical format and granularity, e.g., how many elements (“quantization group”) share a common scale factor. For example, an INT4, per-tensor quantizer first computes the scale as the maximum absolute value of the entire tensor, scales the elements to the maximum representable range of INT4 [-7, +7], and then casts to INT4 with rounding: $\hat{A}=\lceil A/\delta_{A}\rfloor,\delta_{A}=\max(|{A}|)/7$. The dequantization process is an element-wise scaling. For example, for per-tensor dequantization, $\psi_{\delta_{A}\delta_{B}}^{-1}(\hat{A}\hat{B})=\hat{A}\hat{B}\times\delta_{A}\delta_{B}$.
Table 1: Speedup compared to matrix multiplication in FP16 with an FP32 accumulator.
GPU MM Input MM Accumulator Speedup
RTX4090 FP16 FP16 2x
FP8 FP32 2x
L40, L20 H100 FP16 FP16 1x
FP8 FP32 2x
@@ 2.3 SageAttention
Based on the block tiling in FlashAttention (Dao et al., 2022), SageAttention (Zhang et al., 2025) quantizes $Q,K$ to INT8 in a per-block granularity, i.e., each $Q_{i},K_{i}$ has a separate scalar scale: $\delta_{Q_{i}}=\max(|{Q_{i}|})/127,\delta_{K_{j}}=\max(|{K_{j}|})/127$. In this way, the product $S_{ij}$ in Eq. (1) can be approximated as $S_{ij}\approx\hat{Q}_{i}\hat{K}_{j}^{\top}\times(\delta_{Q_{i}}\delta_{K_{j}}/\sqrt{d})$. To maintain the accuracy, SageAttention proposes a preprocessing technique to by subtracting the token-wise mean from $K$. Additionally, SageAttention keeps $\widetilde{P}_{ij}$ and $V_{j}$ in FP16, but utilizes an FP16 accumulator (rather than FP32) for computing the product $\widetilde{P}_{ij}V_{j}$. Reducing the accumulator precision can accelerate the matrix multiplication (MM) on RTX4090 GPU. However, other GPUs, such as L20, L40, or H100, do not exhibit this behavior, as shown in Table 1.
3 SageAttention2
In this section, we introduce SageAttention2, an efficient and accurate quantized attention. The workflow of SageAttention2 is shown in Fig. 2. We quantize $Q,K$ to INT4 and $\tilde{P},V$ to FP8 to maximize the efficiency and propose several techniques, including $QK$-smoothing, per-thread quantization, and two-level accumulation to preserve the accuracy, which we shall discuss in subsequent subsections.